OpenStudy (anonymous):

how do i simplify (1/3) with the exponent of 4

OpenStudy (amistre64):

(1/3)^4 = 1/81

OpenStudy (amistre64):

1/3.3.3.3 = 1/9.9 = 1/81

OpenStudy (anonymous):

-1/64

OpenStudy (amistre64):

umm..... no

OpenStudy (anonymous):

my bad i posted this on the wrong one